OpenStudy (robyon03):

To test the effectiveness of a gunpowder mixture, 1 gram was exploded under controlled STP conditions, and the reaction chamber was found to expand to 310 cm3. What volume would these gases occupy at the temperature produced by an uncontrolled explosion (2,200°C) and 2.1 atmospheres of pressure?

OpenStudy (cuanchi):

P1V1/T1=P2V2/T2 => V2 = (P1V1T2)/(T1P2) convert the T2 to kelvin 2200°C + 273 = ??? P1= 1atm P2= 2.1 atm T1= 273 T2= 2200 + 273= V1= 310cm3